# XAL: EXplainable Active Learning Makes Classifiers Better Low-resource Learners

### v3

@@ 1 Introduction
We conduct experiments on various text classification tasks involving different level of difficulty in understanding rationals. Experimental results manifest that XAL can achieve substantial improvement in all tasks. Ablation studies demonstrate the effectiveness of each component, and human evaluation shows that the model trained in XAL works well in explaining its prediction. XAL also demonstrates superior performance with only 500 instances when compared to in-context learning by ChatGPT, underscoring the effectiveness of our model at a minimal cost. To our knowledge, we are the first to incorporate the model’s explanation (explanation score) to improve the effectiveness of data selection in AL process. We will release the code and data to facilitate further research.
@@ 2 Related Work
Large Language Model. Recently, LLMs of generative schema have shown excellent performance in various NLP tasks. However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), inclination to their internal knowledge Yan et al. (2024), catastrophic forgetting during instruction tuning Luo et al. (2023b), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021). Considering the problems, LLMs are also applied to improve the generalization of smaller models for specific tasks. Some studies distill the knowledge of large language models to a smaller one Hsieh et al. (2023) and some also use LLMs to augment data to achieve stronger text classification performance Ye et al. (2022); Yu et al. (2023). LLMs has also demonstrated a strong capability in generating high-quality reasoning steps Hsieh et al. (2023); Wei et al. (2022); Kojima et al. (2022). In this study we aim to distill the reasoning ability of LLMs to smaller models, and encourage the models to distinguish the reasonablity of explanations to identify informative data in AL scenario.
@@ 4 Experiments
Task
Dataset
# Labels Train Dev Test
Natural Language Inference
RTE (Bentivogli et al., 2009)
2 2,240 250 278
Paraphrase Detection
MRPC (Dolan et al., 2004)
2 3,667 409 1,726
Stance Detection
COVID19 (Glandt et al., 2021)
3 4,533 800 800
Category Sentiment Classification
MAMS (Jiang et al., 2019)
3 7,090 888 901
(Dis)agreement Detection
DEBA (Pougué-Biyong et al., 2021)
3 4,617 578 580
Relevance Classification
CLEF (Kanoulas et al., 2017)
2 7,847 981 982
@@ 4.1 Tasks and Dataset
We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text Bentivogli et al. (2009); Wang et al. (2018); (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff Jiang et al. (2019); Luo et al. (2022a); (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target Glandt et al. (2021); Luo et al. (2022b); (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment Pougué-Biyong et al. (2021); Luo et al. (2023a); (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic Kanoulas et al. (2017); Guo et al. (2023). Appendix A demonstrates the details of six datasets with examples.  [^1] Without losing generality, we randomly split the training set in RTE, and MRPC into train/dev set with proportion 9:1. In DEBA, we adopt the topic of climate change for experiments.
@@ 5.1 Main Results
We mainly consider two different settings: (1) Given the data selection budget, we observe the trend of changes in model performance; (2) Given the performance upper bound, we observe the number of required instances that the model needs to achieve 90% of the upper-bound performance. We utilize FLAN-T5-large Chung et al. (2022) as our backbone network 222 We implement baselines using FLAN-T5-large as well., ChatGPT is adopted to generate the explanations. The implemented details can be found in Appendix D and the detailed values of the result can be found in Appendix E.1.
@@ 5.3 Explanation Generation
We also carry out experiments to analyze how the generation of explanations impacts model performance. Specifically, we replace ChatGPT with ALPACA-7B (Taori et al., 2023), and GPT4  [^5] https://openai.com/gpt-4 to generate explanations on the MAMS dataset. The results are presented in Table 2. We also observe that the ALPACA-7B can also provide useful explanations to some extent and enhance the model performance compared with ME. This suggests that LLMs, when used as an assistant in XAL, can provide consistent explanation generation and enhance model performance. The results also indicate that the model performance can be affected by the explanation generation model and it is applicable to use open-source LLMs. The results of human annotation are also discussed in Appendix F.
@@ Appendix A Tasks and Corresponding Prompts
Task
Text
Natural Language Inference
Sentence 1: Danny Kennedy, Greenpeace campaigns director, said: "The burden of proof in the Scott Parkin expulsion case lies morally with the Commonwealth, to prove that he is a danger. When the Government brought in anti-terror legislation, they promised the public that these laws would only be used to confront a real and present risk of a terrorist attack, not a sweep-all approach against citizens. Peace is not terrorism. Peace is not a threat to national security. No democratic government should expel a foreign citizen because [it] opposes his political opinions."
Sentence 2: Greenpeace director said that peace is terrorism.
Label: Not Entailment.
Paraphrase Detection
Sentence 1: Last week the power station’s US owners, AES Corp, walked away from the plant after banks and bondholders refused to accept its financial restructuring offer .",
Sentence 2: "The news comes after Drax’s American owner, AES Corp. AES.N, last week walked away from the plant after banks and bondholders refused to accept its restructuring offer.
Label: Paraphrase/Semantic Equivalent.
Category Sentiment Classification
Text: I left feeling unsatisfied, except for having a nice chance to people watch in the cozy atmosphere with my over-priced pasta bolognese.
Target: Ambience
Label: Positive
Stance Detection
Text: Michigan is fining individuals 500\$ for not wearing a mask in public. How do y’all feel about this? Curious because I am torn about being so forceful but agree that people should wear masks. #MaskOn.
Target: Face Mask
Label: Favor
(Dis)agreement Detection
Text 1: True, but with lower power usage, you have less heat to dissipate, meaning you can overclock it even more.
Text 2: AMD creates a chip that saves energy by over 31 times. Someone show this to r/PCMasterRace cause we need to switch to AMD.
Label: Agree.
Relevance Classification
Document: 99mtechnetium penicillamine: a renal cortical scanning agent. 99mTechnetium penicillamine, a renal cortical imaging agent, can be used to provide a rapid, safe, and non-invasive assessment of renal morphology and the renal vascular supply. Since this agent is not excreted significantly during the imaging procedure cortical scans of high quality can be obtained without image deterioration owing to a superimposed collecting system. These scans, which are clearly superior in anatomical detail to earlier scans using 131I hippuran, can be obtained along with the 131I hippuran renogram when the patient comes to the nuclear medicine department. Herein we demonstrate the anatomical detail it is now possible to achieve by presenting the cortical renal scans and accompanying radiograms from 5 patients with different renal pathology.
Topic: Procalcitonin, C-reactive protein, and erythrocyte sedimentation rate for the diagnosis of acute pyelonephritis in children.
Label: Not Relevant.
@@
Task
Prompts
Label set
Natural Language Inference
Sentence 1: {Text 1}. Sentence 2: {Text 2}. Sentence 1 can {Label} sentence 2, explain the reason within 50 words.
{Entail, Not Entail}
Paraphrase Detection
Sentence 1: {Text 1}. Sentence 2: {Text 2}. The relation between the above two sentences is {Label}, explain the reason within 50 words.
{Paraphrase Equivalent, Not Paraphrase Equivalent}
Category Sentiment Classification
{Text}.
{Positive, Negative, Neutral}
Stance Detection
{Text}. The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words.
{Favor, Neutral, Against}
(Dis)agreement Detection
Sentence 1: {Text 1}. Sentence 2: {Text 2}. The stance of sentence 2 is {Label} to sentence 1, explain the reason within 50 words.
{Agree, Neutral, Disagree}
Relevance Classification
{Document}. The scientific document is {Label}’ to the research topic {Topic}, explain the reason within 100 words.
{Relevant, Not Relevant}
@@ Appendix B Explanation Examples
Input
Label
Explanation
RTE
Sentence 1: The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes.
Sentence 2: Nobel Peace Prize candidates have been chosen.
Not Entailment
@@
Sentence 1 states that the Nobel prizes are the main forms of recognition for various fields, including peace. However, sentence 2 only mentions that candidates have been chosen for the Nobel Peace Prize, without specifying if they have actually received the prize. Therefore, sentence 2 does not necessarily follow from sentence 1.
Entailment
@@
Sentence 1 can entail sentence 2 because sentence 1 mentions that one of the main forms of recognition for those working for peace is the Nobel prizes. Therefore, it is reasonable to assume that candidates for the Nobel Peace Prize have been chosen.
MRPC
Sentence 1: Maddox, 87, cracked two ribs when he fell about 10 days ago at an assisted living home where he was recovering from intestinal surgery, Virginia Carnes said.
Sentence 2: Maddox, who had battled cancer since 1983, cracked two ribs earlier this month when he fell at an assisted living home where he was recovering from surgery.
Equivalent
@@
The two sentences provide similar information but are expressed differently. Sentence 1 mentions that Maddox fell about 10 days ago, while Sentence 2 specifies that the fall occurred earlier this month. Despite the slight difference in wording, both sentences convey the same meaning.
Not Equivalent
@@
The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information about the timing and reason for Maddox’s rib injury. In the first sentence, it mentions that Maddox fell about 10 days ago while recovering from intestinal surgery, whereas the second sentence states that he fell earlier this month while recovering from surgery without specifying the exact time frame.
COVID19
Tweet: Placed2nd @AngelaBelcamino Omg people say I have to wear a mask and I’m crying because I have to wear a mask to save lives boo hoo!! MAGAts are stupid!!!
Favor
@@
The stance of this tweet towards face masks is favorable. The reason for this is that the tweet expresses support for wearing masks to save lives and criticizes those who oppose wearing masks, referring to them as ""MAGAts"" (a term often used to describe supporters of former President Donald Trump)."
Against
@@
The reason behind the stance against face masks in this tweet seems to be frustration and ridicule towards those who oppose wearing masks. The tweet suggests that the person finds the idea of wearing masks to save lives unnecessary and mocks those who support it, particularly referring to them as ""MAGAts,"" a derogatory term for supporters of former President Trump.
Neutral
@@
The tweet expresses a neutral stance towards face masks by mocking the idea of crying over having to wear one. While it criticizes MAGA supporters, it does not explicitly state a negative or positive opinion about face masks themselves."
@@ I.1 Model Generation
Input
Golden
Explanation
i. RTE
Sentence 1: Dana Reeve, the widow of the actor Christopher Reeve, has died of lung cancer at age 44, according to the Christopher Reeve Foundation.
Sentence 2: Christopher Reeve had an accident.
Not Entailment
Sentence 1 states that Dana Reeve, the widow of Christopher Reeve, has died of lung cancer at the age of 44. It does not provide any information about what caused Christopher Reeve’s accident. Therefore, sentence 1 does not entail sentence 2.
ii. MRPC
Sentence 1: Air Commodore Quaife said the Hornets remained on three-minute alert throughout the operation.
Sentence 2: Air Commodore John Quaife said the security operation was unprecedented.
Not Equivalent
The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information. Sentence 1 states that the Hornets were on three-minute alert during the operation, while Sentence 2 states that the security operation was unprecedented.
iii. MAMS
Review: Waiter took our drink order and then we didn’t see him for 15 minutes.
Category: Stuff.
Negative
The reason for the negative sentiment towards the staff in this review is because the waiter took the drink order and disappeared after 15 minutes. This indicates a lack of professionalism and attentiveness, which can be frustrating for customers.
iv. COVID19
Tweet: @WCheesemaker @StephenKing No. You’re missing the point. And from the way I phrased my statement I can understand why. I was attempting to point out masks work! They help stop the spread of disease. I should have stated it more clearly. I’m a scientist; a big fan of masks- and science!! :relaxed:
Target: Face mask.
Favor
The stance of this tweet is in favor of face masks. The reason given is that the tweeter believes masks are effective in stopping the spread of disease, and therefore should be encouraged to use them.
v. DEBA
Sentence 1: Best thing you can do is simply not drive. Take public transportation instead.
Sentence 2: That’s hugely unrealistic for many people living in places with poor public transport provisions (rural areas, poorer areas) or who live in poorly planned towns where amenities are spatially distant.
Disagree
Sentence 2 disagrees with Sentence 1 because it argues that taking public transportation instead of driving is not feasible for many people living in places with poor public transport provisions or in poorly planned towns.
vi. DEBA
Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters. This is… uh, not a guy who’s gonna accomplish much in office.
Sentence 2: Biden has surprised me with some of his action but on the whole it’s still up to activists and local politicians to get stuff done. Shame
Agree
Sentence 2 agrees with Sentence 1 because it acknowledges that Biden has made some progress, but it also emphasizes that it is up to activists and local politicians to bring about meaningful change.

### v4

@@ Abstract
Yun Luo1,3    Zhen Yang2    Fandong Meng2    Yingjie Li 1    Fang Guo1,3
@@
1School of Engineering, Westlake University   2WeChat AI, Tencent Inc.
3 Zhejiang University 4 School of Cyber Science and Engineering, Sichuan University
5 Institute of Advanced Technology, Westlake Institute for Advanced Study
{luoyun,zhangyue}@westlake.edu.cn
@@ 1 Introduction
We conduct experiments on various text classification tasks involving different level of difficulty in understanding rationals. Experimental results manifest that XAL can achieve substantial improvement in all tasks. Ablation studies demonstrate the effectiveness of each component, and human evaluation shows that the model trained in XAL works well in explaining its prediction. XAL also demonstrates superior performance with only 500 instances when compared to in-context learning by ChatGPT, underscoring the effectiveness of our model at a minimal cost. To our knowledge, we are the first to incorporate the model’s explanation (explanation score) to improve the effectiveness of data selection in AL process. The codes and data have been released in the link to facilitate further research  [^1] https://github.com/LuoXiaoHeics/XAL.
@@ 2 Related Work
Large Language Model. Recently, LLMs of generative schema have shown excellent performance in various NLP tasks bubeck2023sparks,yang2023teal. However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), inclination to their internal knowledge Yan et al. (2024), catastrophic forgetting during instruction tuning Luo et al. (2023), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021). Considering the problems, LLMs are also applied to improve the generalization of smaller models for specific tasks. Some studies distill the knowledge of large language models to a smaller one Hsieh et al. (2023) and some also use LLMs to augment data to achieve stronger text classification performance Ye et al. (2022); Yu et al. (2023). LLMs has also demonstrated a strong capability in generating high-quality reasoning steps Hsieh et al. (2023); Wei et al. (2022); Kojima et al. (2022). In this study we aim to distill the reasoning ability of LLMs to smaller models, and encourage the models to distinguish the reasonablity of explanations to identify informative data in AL scenario.
@@ 4 Experiments
Task Dataset # Labels Train Dev Test
Natural Language Inference RTE (Bentivogli et al., 2009) 2 2,240 250 278
Paraphrase Detection MRPC (Dolan et al., 2004) 2 3,667 409 1,726
Stance Detection COVID19 (Glandt et al., 2021) 3 4,533 800 800
Category Sentiment Classification MAMS (Jiang et al., 2019) 3 7,090 888 901
(Dis)agreement Detection DEBA (Pougué-Biyong et al., 2021) 3 4,617 578 580
Relevance Classification CLEF (Kanoulas et al., 2017) 2 7,847 981 982
@@ 4.1 Tasks and Dataset
We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text; (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff; (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target; (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment; (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic. The details of the dataset we used are shown in Table 1. Appendix A demonstrates the details and prompts of six datasets with examples.  [^2] Without losing generality, we randomly split the training set in RTE, and MRPC into train/dev set with proportion 9:1. In DEBA, we adopt the topic of climate change for experiments.
@@ 5.1 Main Results
We mainly consider two different settings: (1) Given the data selection budget, we observe the trend of changes in model performance; (2) Given the performance upper bound, we observe the number of required instances that the model needs to achieve 90% of the upper-bound performance. We utilize FLAN-T5-large Chung et al. (2022) as our backbone network 333 We implement baselines using FLAN-T5-large as well., ChatGPT is adopted to generate the explanations. The implemented details can be found in Appendix D and the detailed values of the result can be found in Appendix E.1.
@@ 5.3 Explanation Generation
We also carry out experiments to analyze how the generation of explanations impacts model performance. Specifically, we replace ChatGPT with ALPACA-7B (Taori et al., 2023), and GPT4  [^6] https://openai.com/gpt-4 to generate explanations on the MAMS dataset. The results are presented in Table 2. We also observe that the ALPACA-7B can also provide useful explanations to some extent and enhance the model performance compared with ME through our framework. This suggests that LLMs, when used as an assistant in XAL, can provide consistent explanation generation and enhance model performance. The results also indicate that the model performance can be affected by the explanation generation model and it is applicable to use open-source LLMs. The results of human annotation are also discussed in Appendix F.
@@ 8 Ethical Statement
9 Acknowledgement
We acknowledge financial support of the National Natural Science Foundation of China Key Program under Grant Number 62336006.
@@ Appendix A Tasks and Corresponding Prompts
Task Text
Natural Language Inference Sentence 1: Danny Kennedy, Greenpeace campaigns director, said: "The burden of proof in the Scott Parkin expulsion case lies morally with the Commonwealth, to prove that he is a danger. When the Government brought in anti-terror legislation, they promised the public that these laws would only be used to confront a real and present risk of a terrorist attack, not a sweep-all approach against citizens. Peace is not terrorism. Peace is not a threat to national security. No democratic government should expel a foreign citizen because [it] opposes his political opinions." Sentence 2: Greenpeace director said that peace is terrorism. Label: Not Entailment.
Paraphrase Detection Sentence 1: Last week the power station’s US owners, AES Corp, walked away from the plant after banks and bondholders refused to accept its financial restructuring offer .", Sentence 2: "The news comes after Drax’s American owner, AES Corp. AES.N, last week walked away from the plant after banks and bondholders refused to accept its restructuring offer. Label: Paraphrase/Semantic Equivalent.
Category Sentiment Classification Text: I left feeling unsatisfied, except for having a nice chance to people watch in the cozy atmosphere with my over-priced pasta bolognese. Target: Ambience Label: Positive
Stance Detection Text: Michigan is fining individuals 500\$ for not wearing a mask in public. How do y’all feel about this? Curious because I am torn about being so forceful but agree that people should wear masks. #MaskOn. Target: Face Mask Label: Favor
(Dis)agreement Detection Text 1: True, but with lower power usage, you have less heat to dissipate, meaning you can overclock it even more. Text 2: AMD creates a chip that saves energy by over 31 times. Someone show this to r/PCMasterRace cause we need to switch to AMD. Label: Agree.
Relevance Classification Document: 99mtechnetium penicillamine: a renal cortical scanning agent. 99mTechnetium penicillamine, a renal cortical imaging agent, can be used to provide a rapid, safe, and non-invasive assessment of renal morphology and the renal vascular supply. Since this agent is not excreted significantly during the imaging procedure cortical scans of high quality can be obtained without image deterioration owing to a superimposed collecting system. These scans, which are clearly superior in anatomical detail to earlier scans using 131I hippuran, can be obtained along with the 131I hippuran renogram when the patient comes to the nuclear medicine department. Herein we demonstrate the anatomical detail it is now possible to achieve by presenting the cortical renal scans and accompanying radiograms from 5 patients with different renal pathology. Topic: Procalcitonin, C-reactive protein, and erythrocyte sedimentation rate for the diagnosis of acute pyelonephritis in children. Label: Not Relevant.
@@
Task Prompts Label set
Natural Language Inference Sentence 1: {Text 1}. Sentence 2: {Text 2}. Sentence 1 can {Label} sentence 2, explain the reason within 50 words. {Entail, Not Entail}
Paraphrase Detection Sentence 1: {Text 1}. Sentence 2: {Text 2}. The relation between the above two sentences is {Label}, explain the reason within 50 words. {Paraphrase Equivalent, Not Paraphrase Equivalent}
Category Sentiment Classification {Text}. {Positive, Negative, Neutral}
Stance Detection {Text}. The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words. {Favor, Neutral, Against}
(Dis)agreement Detection Sentence 1: {Text 1}. Sentence 2: {Text 2}. The stance of sentence 2 is {Label} to sentence 1, explain the reason within 50 words. {Agree, Neutral, Disagree}
Relevance Classification {Document}. The scientific document is {Label}’ to the research topic {Topic}, explain the reason within 100 words. {Relevant, Not Relevant}
@@ Appendix B Explanation Examples
Input Label Explanation
@@
RTE Sentence 1: The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes. Sentence 2: Nobel Peace Prize candidates have been chosen. Not Entailment Sentence 1 states that the Nobel prizes are the main forms of recognition for various fields, including peace. However, sentence 2 only mentions that candidates have been chosen for the Nobel Peace Prize, without specifying if they have actually received the prize. Therefore, sentence 2 does not necessarily follow from sentence 1.
@@
Entailment Sentence 1 can entail sentence 2 because sentence 1 mentions that one of the main forms of recognition for those working for peace is the Nobel prizes. Therefore, it is reasonable to assume that candidates for the Nobel Peace Prize have been chosen.
@@
MRPC Sentence 1: Maddox, 87, cracked two ribs when he fell about 10 days ago at an assisted living home where he was recovering from intestinal surgery, Virginia Carnes said. Sentence 2: Maddox, who had battled cancer since 1983, cracked two ribs earlier this month when he fell at an assisted living home where he was recovering from surgery. Equivalent The two sentences provide similar information but are expressed differently. Sentence 1 mentions that Maddox fell about 10 days ago, while Sentence 2 specifies that the fall occurred earlier this month. Despite the slight difference in wording, both sentences convey the same meaning.
@@
Not Equivalent The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information about the timing and reason for Maddox’s rib injury. In the first sentence, it mentions that Maddox fell about 10 days ago while recovering from intestinal surgery, whereas the second sentence states that he fell earlier this month while recovering from surgery without specifying the exact time frame.
@@
COVID19 Tweet: Placed2nd @AngelaBelcamino Omg people say I have to wear a mask and I’m crying because I have to wear a mask to save lives boo hoo!! MAGAts are stupid!!! Favor The stance of this tweet towards face masks is favorable. The reason for this is that the tweet expresses support for wearing masks to save lives and criticizes those who oppose wearing masks, referring to them as ""MAGAts"" (a term often used to describe supporters of former President Donald Trump)."
@@
Against The reason behind the stance against face masks in this tweet seems to be frustration and ridicule towards those who oppose wearing masks. The tweet suggests that the person finds the idea of wearing masks to save lives unnecessary and mocks those who support it, particularly referring to them as ""MAGAts,"" a derogatory term for supporters of former President Trump.
@@
Neutral The tweet expresses a neutral stance towards face masks by mocking the idea of crying over having to wear one. While it criticizes MAGA supporters, it does not explicitly state a negative or positive opinion about face masks themselves."
@@ I.1 Model Generation
Input Golden Explanation
i. RTE Sentence 1: Dana Reeve, the widow of the actor Christopher Reeve, has died of lung cancer at age 44, according to the Christopher Reeve Foundation. Sentence 2: Christopher Reeve had an accident. Not Entailment Sentence 1 states that Dana Reeve, the widow of Christopher Reeve, has died of lung cancer at the age of 44. It does not provide any information about what caused Christopher Reeve’s accident. Therefore, sentence 1 does not entail sentence 2.
ii. MRPC Sentence 1: Air Commodore Quaife said the Hornets remained on three-minute alert throughout the operation. Sentence 2: Air Commodore John Quaife said the security operation was unprecedented. Not Equivalent The reason why the two sentences are not paraphrases or semantic equivalents is because they provide different information. Sentence 1 states that the Hornets were on three-minute alert during the operation, while Sentence 2 states that the security operation was unprecedented.
iii. MAMS Review: Waiter took our drink order and then we didn’t see him for 15 minutes. Category: Stuff. Negative The reason for the negative sentiment towards the staff in this review is because the waiter took the drink order and disappeared after 15 minutes. This indicates a lack of professionalism and attentiveness, which can be frustrating for customers.
iv. COVID19 Tweet: @WCheesemaker @StephenKing No. You’re missing the point. And from the way I phrased my statement I can understand why. I was attempting to point out masks work! They help stop the spread of disease. I should have stated it more clearly. I’m a scientist; a big fan of masks- and science!! :relaxed: Target: Face mask. Favor The stance of this tweet is in favor of face masks. The reason given is that the tweeter believes masks are effective in stopping the spread of disease, and therefore should be encouraged to use them.
v. DEBA Sentence 1: Best thing you can do is simply not drive. Take public transportation instead. Sentence 2: That’s hugely unrealistic for many people living in places with poor public transport provisions (rural areas, poorer areas) or who live in poorly planned towns where amenities are spatially distant. Disagree Sentence 2 disagrees with Sentence 1 because it argues that taking public transportation instead of driving is not feasible for many people living in places with poor public transport provisions or in poorly planned towns.
vi. DEBA Sentence 1: I can’t imagine believing Joe "Nothing Will Fundamentally Change" Biden is going to do anything except continue to serve his corporate masters. This is… uh, not a guy who’s gonna accomplish much in office. Sentence 2: Biden has surprised me with some of his action but on the whole it’s still up to activists and local politicians to get stuff done. Shame Agree Sentence 2 agrees with Sentence 1 because it acknowledges that Biden has made some progress, but it also emphasizes that it is up to activists and local politicians to bring about meaningful change.